What is the length of a column of air in a closed pipe if it vibrates with a fundamental mode of vibration of 160Hz? B) What would be the next shortest length? Assuming the speed of sound to be 340m/sec.
1
Expert's answer
2018-01-24T09:35:08-0500
A)
f=v/4L
L=v/4f=340/4(160) =0.531 m.
B)
L^'=3L=3( 340/4(160) )= 1.594 m.